OpenStudy (anonymous):

what is the slope of the line that contains the points (-2, 7) and (2, 3)

OpenStudy (anonymous):

Do difference of y/ difference of x

OpenStudy (anonymous):

So the difference in y can be 7-3=5

OpenStudy (marihelenh):

\[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }=m\]

OpenStudy (shamim):

Slope m=(y2-y1)/(x2-x1)

OpenStudy (marihelenh):

m=slope

OpenStudy (anonymous):

Could also be y1-y2/X1-x2

OpenStudy (shamim):

Given x1=-2 y1=7 x2=2 y2=3

OpenStudy (anonymous):

It equals the same thing

OpenStudy (shamim):

Welcome to openstudy @smartguy122

OpenStudy (anonymous):

the choices is A.-4 B.-1 c.4 D.1

OpenStudy (marihelenh):

\[\frac{ 7-3 }{ -2-2 }=\frac{ 4 }{ -4 }=-1\]

OpenStudy (shamim):

None will tell u the result directly

OpenStudy (marihelenh):

whoops

OpenStudy (shamim):

U need to understand the procedure of doing ur math problem

OpenStudy (anonymous):

so is it -1?

OpenStudy (marihelenh):

Yeah

OpenStudy (anonymous):

thanks